

********************************************************************
UARS WEEKLY STATUS REPORT - 30 APRIL 1999
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
After turning back on April 11th, SUSIM gathered its normal solar,
offset, and occultation data without apparent problems.
Reprocessing of the SUSIM data to version 20 is now complete through
UARS day 2523.
***********
PEM
***********
PEM is active under PMON control. Energy processor #6 for the HEPS-1 unit
is activating correctly now. The PEM team and FOT found that the entanglement
of energy processor #6 was related to the differences between OBC activation
commanding and activation commanding from the ground. Attempts at adjusting OBC
command timing failed until the correct combination of command sequence was
found. The FOT and the PEM team are completing close-out documentation on this
problem.
While under PMON control, the PEM instrument computes energy deposition
from the particle spectrometers.
***********
HALOE
***********
HALOE was operating successfully prior to its scheduled turn off on April
27, 1999. All engineering parameters were nominal and the science data
results are of excellent quality.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
WINDII successfully completed its observation period of April 25-29.
Reprocessing and the evaluation of reprocessed data are continuing.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
Y2K end-to-end test successfully completed Thursday April 29, 1999.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
1 day CLA (7.0) PROD (TEST) (L1-L3) (1R-1C=0L)
50 days PEM (3.32) PROD (OA-L2) (760R-578C-1F=181L)
1 day SOL (4.62) TEST (DEBUG)(L2-L3) (1R-1F=0L)
111 days WIN (5.11) PROD (LOG) (2014R-1303C=711L)
76 days WIN (5.11) PROD (L1-L3) (2014R-339-3F=1672L).
DAAC data created for February 1999 has been loaded into the DAAC transfer
tables. DAAC data thru January 1999 for HALOE, HRDI, PEM, NMC, and UKMO is
revealed/available to the User community. The CDHF will load March 1999
current day DAAC data in May 1999.
SOLSTICE "The debug SOLSTICE job was resubmitted with debug changes, and again
aborted due to "general unrecoverable error", this time reported by
ASGCAL. Investigation of this error will take place next week."
CDHF pending software upgrades: DEC AXP Fortran 6.5, DEC AXP OPEN
VMS 6.2, OpenIngres.
***********
FOT
***********
This report covers 17 April 1999 (Orbit 41515, GMT Day 107 through
23 April GMT Day 113).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HRDI 107/0037 41515 Day Only
PEM 107/0040 41515 AXIS On
WINDII 109/2307 41559 Off
PEM 109/2307 41559 AXIS-2 Off
HRDI 113/0059 41605 Full Time
The beta angle ranged between the angles of 46.5 degrees to 30.0 degrees
for this report period. The beta angle is now decreasing toward a minimum
angle of 2.9 degrees on 30 April 1999 (DOY 120).
Two manually generated Stellar offset maneuvers were performed on GMT days
111 and 112 at the request of SOLSTICE operations. The offsets are similar
to the normal weekly SOLSTICE offset which usually occurs during a Solar
tracking session.
A series of SSPP gridded offset maneuvers were performed on GMT days 107
through 112 at the request of SUSIM operations. The offsets utilized the 5
and 10 arc-minute interval steps, as opposed to the optional 15 and 30 arc-
minute intervals.
A UARS TDRS event was scheduled partially out of view on GMT day 112 due to
FOT error. This event was not one of UARS' originally scheduled events,
but was added in real time to replace an event that NCC had taken away for
a higher priority TDRS user. While scheduling the new event, the FOT
mistakenly used only the TDRS view periods contained on the UPS graphical
display, instead of the CMS generated HG Antenna Tracking report. Due to
slewing limitations, the view periods on HG tracking report are often more
restrictive than the view periods on the UPS, because the UPS does not take
into account the physical operation of the HG antenna. In this particular
case, the High Gain Antenna began to slew to the other TDRS approximately 3
minutes into the newly scheduled event, resulting in a tape recorder
playback data loss. An Anomaly Investigation Report (AIR 99-022) is being
written to document this problem.
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 1.0 to 1.3 degrees C. Additional battery performance data
for each battery on 16 April 1999 and 23 April 1999 are provided in
Appendix D.
The clock error ranged from -11.1 to 10.8 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 03
% HRDI 01
% PEM 02
% WINDII 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 05
Five (5) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 99/054 - 2287.496923
Transponder B: 99/055 - 2287.498750
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) was not
available this week.
The FOT is still awaiting the delivery of the 8-week Ground Track Y2K test
file from FDF. This file was expected to be delivered this week, but was
not received as of the end of this reporting period.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
FOT personnel are supporting the IMOC (SCS-21) development effort. The
current work is a continuation of the SCS-21 UARS implementation effort
done last year which culminated in a level 1 demonstration.
A first cut of the high level UARS mission support requirements for the
IMOC were completed on Friday. These requirements were presented and
discussed in a Friday Meeting with Evan Eller.
The FOT continues to investigate the feasibility of a new PMON action to
change battery V/T curves based on loading to mitigate battery risks in sun
pointing safehold.
YEAR 2000 OPERABILITY SUPPORT - No change
UARS PROPULSION MODULE B/D THRUSTER LINE FILL - No change
MEETINGS
FOT supported the weekly IMOC meeting on 21 April 1999.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver (reverse-to-forward) is scheduled for 30
April 1999 (DOY 120).
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 23 April 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
41515 107/0040 99-140 PEM AXIS Power ON
41517 107/0402 99-141 OBC Clk Adj fr 20.6 to 0.00 msec/day
41547 109/0404 99-142 OBC Clk Adj fr 0.00 to 20.6 msec/day
41559 109/2307 99-143 WINDII Power ON
41559 109/2307 99-144 AXIS2 Power OFF
41570 110/1701 99-145 HRDI PMON Exercise
41575 111/0059 99-146 OBC Clk Adj fr 20.6 to 0.00 msec/day
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 23 April 1999
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 23 April 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
41517 107/04:04 (17 Apr 99) from 20.6 to 00.0 msec/day
41547 109/04:10 (19 Apr 99) from 00.0 to 20.6 msec/day
41575 111/01:02 (21 Apr 99) from 20.6 to 00.0 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 23 April 1999
16 Apr 1999 - GMT Day 106 Beta = 46.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -56.0/-78.4 -3.05/-4.31 +0.0/ +0.0 16.6 0.000 0.0
2 +78.4/-5.6 +6.44/+4.51 +27.2/-16.0 24.3 1.019 22.1
3 +33.6/-33.6 +5.15/+3.87 +27.2/-16.8 24.2 1.023 21.8
23 Apr 1999 - GMT Day 113 Beta = 30.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -67.2/-89.6 -3.05/-4.31 +0.0/ +0.0 16.6 0.000 0.0
2 +56.0/-5.6 +6.11/+4.51 +30.4/-15.2 24.8 1.021 22.0
3 +44.8/-11.2 +5.15/+3.23 +29.6/-15.6 24.8 1.024 21.7
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 23 April 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-019 FOT 41573 110/2154 20772 NASCOM PDP Failure
99-020 HRDI 41560 110/0051 N/A HRDI
99-016 FOT 41588 111/2157 14170 Generic Late Acquisition
99-021 HRDI 41589 112/0058 N/A HRDI IN PMON MODE
99-022 FOT 41603 112/2149 N/A Data Loss
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-016 FOT see list Apr 99 14170 Generic Late Acq
99-019 FOT 41573 110/2154 20772 NASCOM PDP Failure
99-020 HRDI 41560 110/0051 N/A HRDI
99-021 HRDI 41589 112/0058 N/A HRDI IN PMON MODE
99-022 FOT 41603 112/2149 N/A Data Loss
SEVNTFW99/17